A thermos bottle containing 250g of Coffee at 90°C is added with a 20g of milk at 5°C. After thorough mixing, what is the final temperature. Capacity for water, coffee and milk is 1.00Cal/g°C
1
Expert's answer
2019-04-18T11:33:04-0400

From the conservation of energy:


"Q_c=Q_m"

"m_c c (T_c-T)=m_m c (T-T_m)"

"m_c (T_c-T)=m_m (T-T_m)"

"250 (90-T)=20 (T-5)"

"T=83.7\u00b0C"